OpenStudy (anonymous):

simplify radical 32??

OpenStudy (hawkfalcon):

What is a small(ish) number that can be rooted that fits into that? Like 9,16,25,36.

OpenStudy (anonymous):

2 @hawkfalcon

OpenStudy (hawkfalcon):

Not 2, that can't be rooted. In this case you would use 16.

OpenStudy (hawkfalcon):

For simplifying radicals, find a rootable number that fits into your radical an even amount of times. So, 16.

OpenStudy (hawkfalcon):

Then, 32/16 = 2, so you are left with a 2 in the radical, and the root of 16 outside of it.

OpenStudy (anonymous):

so the answer is 16 radical 2

OpenStudy (hawkfalcon):

No, you have to root the 16, as you moved it outside of the radical. Think of it as \[\sqrt{16} * \sqrt{2} at this point. so \it would be 4\sqrt{2}\]

OpenStudy (hawkfalcon):

Does that make sense?

OpenStudy (anonymous):

yes thanks